OpenStudy (anonymous):

Solve the problem. Use π= 3.14 when necessary. Find the volume of a rectangular pyramid with base 17 m2 and height 7 m. Find the result to the nearest unit. A. 38 m3 B. 119 m3 C. 60 m3 D. 40 m3

OpenStudy (anonymous):

I know that (c ) isnt the answer

OpenStudy (anonymous):

did u get the solution in the previous question?

OpenStudy (anonymous):

yes but this one i dnt get

OpenStudy (anonymous):

volume of pyramid= 1/3 * base* height answer= 40 (approx)

OpenStudy (anonymous):

I did 17x7=119/3

OpenStudy (anonymous):

yes, 119/3= 40 approx no problem